show step by step:

(3x2 + x - 4) - ( 4x - 6 + 8x2)


1
Expert's answer
2021-04-15T07:21:33-0400

"To \ find \ the \ difference \ of \ the \ polynomials \\
(3x^2 +x-4)-(4x-6+8x^2)\\
Opening \ the \ parenthesis \ we \ get \\
3x^2 +x-4 -4x-(-6)-(8x^2 ) \because -a(b+c)=-ab-ac\\"

"3x^2 +x-4-4x+6-8x^2\because (-1)(ax)=-ax\\
Regrouping \ the \ like \ powers\ of \ x\ we \ get \\
3x^2-8x^2+x-4x-4+6=(3-8)x^2+(1-4)x+(-4+6) \\ \because (a+b)x=ax+bx\\
-5x^2 -3x+2\\
\therefore(3x^2+x-4)-(4x-6+8x^2)=-5x^2-3x+2"